A ball of mass 60kg is moving due south with a speed of 50m per second at latitude 30°N.Calculate the magnitude and direction of the coriolis force on the ball.Compare the magnitude of this force to the weight of the ball?
1
Expert's answer
2018-04-29T14:02:08-0400
F_c=2mωv sin⁡α=2(60)(7.3∙〖10〗^(-5) )(50) sin⁡30=0.22 N.
The direction is to the left of motion in the Southern Hemisphere: toward east.
F_c/F_g =(2mωv sin⁡α)/mg=(2ωv sin⁡α)/g=(2(7.3∙〖10〗^(-5) )(50) sin⁡30)/9.8=3.7∙〖10〗^(-4)